# Conditional Distribution Compression via the Kernel Conditional Mean Embedding

### v3

@@ B.11.1 Joint Kernel Inducing Points
Defining $\tilde{\bm{X}}:=[\tilde{\bm{x}}_{1},\tilde{\bm{x}}_{2},\dots,\tilde{\bm{x}}_{m}]^{T}$ and $\tilde{\bm{Y}}:=[\tilde{\bm{y}}_{1},\tilde{\bm{y}}_{2},\dots,\tilde{\bm{y}}_{m}]^{T}$, in JKIP, we optimise the following objective
@@
$=\frac{2}{m^{2}}\tilde{\bm{K}}_{m}(\tilde{\bm{x}}_{t})^{T}\tilde{\bm{L}}_{m}(\tilde{\bm{y}}_{t})-\frac{2}{mn}\bm{K}_{n}(\tilde{\bm{x}}_{t})^{T}\bm{L}_{n}(\tilde{\bm{y}}_{t})$ (31)
where $\tilde{\bm{K}}_{m}(\bm{x}):=[k(\bm{x},\tilde{\bm{x}}_{1}),\dots,k(\bm{x},\tilde{\bm{x}}_{m})]^{T}$, and $\bm{K}_{n}(\bm{x}):=[k(\bm{x},\bm{x}_{1}),\dots,k(\bm{x},\bm{x}_{n})]^{T}$, $\tilde{\bm{L}}_{m}(\bm{y}):=[l(\bm{y},\tilde{\bm{y}}_{1}),\dots,l(\bm{y},\tilde{\bm{y}}_{m})]$, and $\bm{L}_{n}(\bm{y}):=[l(\bm{y},\bm{y}_{1}),\dots,l(\bm{y},\bm{y}_{n})]^{T}$. This reduces the cost of evaluating the objective function by a factor of $m$, both in storage and time.
@@ B.11.2 Average Conditional Kernel Inducing Points
$\underset{\text{(b)}}{=}\frac{2}{n}\tilde{\bm{L}}_{m}(\tilde{\bm{y}}_{t})^{T}W_{\tilde{\bm{X}}\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}K_{\bm{X}\tilde{\bm{X}}}\bm{w}_{t}-\frac{2}{n}\bm{L}_{m}(\tilde{\bm{y}}_{t})^{T}K_{\bm{X}\tilde{\bm{X}}}\bm{w}_{t}$ (32)
where (a) follows from the symmetry of the kernel functions and a simple reordering of the terms, and (b) follows from defining $\bm{w}_{t}$ to be the $t^{\text{th}}$ row of $W_{\tilde{\bm{X}}\tilde{\bm{X}}}$.
@@ Appendix C Experiment Details
$k(\bm{x},\bm{x}^{\prime}):=\exp\left(-\frac{1}{2\alpha_{k}^{2}}\|\bm{x}-\bm{x}^{\prime}\|^{2}\right),\;\;l(y,y^{\prime}):=\exp\left(-\frac{1}{2\alpha_{l}^{2}}\|\bm{y}-\bm{y}^{\prime}\|^{2}\right)$
@@ C.1.2 Matching the Empirical Conditional Distribution - Continuous / Regression
Figure 16: Results for Heteroscedastic data with parameters set as $\bm{a}:=[3,-3,6,-6]^{T}$, $\bm{b}:=[1,0.1,2,0.5]^{T}$, $\bm{c}:=[-5,-2,2,5]^{T}$, $\sigma_{1}^{2}=0.75$, and $\sigma_{2}^{2}=0.1$. RMSE is reported across a variety of test functions, as the size of the compressed set increases. For JKH (orange), JKIP (red), ACKH (blue), and ACKIP (green), we display the median performance (bold line) with the 25th-75th percentiles (shaded region) over 20 runs. The error of random sampling (black) over 500 runs is also plotted for comparison, as well as the performance of the full data (purple).
@@ C.1.3 Matching the Empirical Conditional Distribution - Discrete / Classification
Synthetic: We generate an unbalanced 4-class dataset using the multinomial logistic regression model, where conditional class probabilities are given by $\mathbb{P}(Y=0\mid X=\bm{x})=\frac{1}{1+\sum_{j=1}^{3}\exp({\bm{\beta}_{j}\cdot\bm{x}})}$ and $\mathbb{P}(Y=k\mid X=\bm{x})=\frac{\exp({\bm{\beta}_{k}\cdot\bm{x}})}{1+\sum_{j=1}^{3}\exp({\bm{\beta}_{j}\cdot\bm{x}})}$, $1\leq k\leq 3$, and $\mathbb{P}_{X}$ is a 2D Gaussian mixture model with 100 components. We assign classes using $\bm{\beta}:=\begin{bmatrix}10&8&1&45\\ 40&45&40&10\end{bmatrix}^{T}$ and, to ensure class overlap, introduce additive noise $\epsilon\sim\mathcal{N}(0,100)$ to the exponential. This setup results in class proportions of approximately 32% (class 0), 12% (class 1), 19% (class 2), and 37% (class 3).
@@ C.3.1 Deriving the Marginal Expectation
$\mathbb{E}_{\bm{x}\sim\mathbb{P}_{X}}\left[k(\bm{x},\bm{x}^{\prime})k(\bm{x},\bm{x}^{\prime\prime})\right]=\int_{\mathcal{X}}k(\bm{x},\bm{x}^{\prime})k(\bm{x},\bm{x}^{\prime\prime})\text{d}\bm{x}$
@@ C.3.2 Deriving the Joint Expectation
$\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}A\bm{\omega}+\bm{b}^{T}\bm{\omega}+c\right)\text{d}\bm{\omega}$
for $\bm{\omega}:=(\bm{x},\bm{y})^{T}$ as by completing the square, it can be shown [86] that
$\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{T}A\bm{\omega}+\bm{b}^{T}\bm{\omega}+c\right)\text{d}\bm{\omega}=\frac{2\pi}{|A|^{\frac{1}{2}}}\exp\left(c+\frac{1}{2}\bm{b}^{T}A^{-1}\bm{b}\right).$
@@
$=\exp\left(-\frac{1}{2\alpha_{k}^{2}\alpha_{l}^{2}}\left[\begin{pmatrix}\bm{x}\\ \bm{y}\end{pmatrix}^{T}\begin{pmatrix}\alpha_{l}^{2}&0\\ 0&\alpha_{k}^{2}\end{pmatrix}\begin{pmatrix}\bm{x}\\ \bm{y}\end{pmatrix}+\begin{pmatrix}-2\lambda^{2}_{l}\bm{x}^{\prime}\\ -2\alpha_{k}^{2}\bm{y}^{\prime}\end{pmatrix}^{T}\begin{pmatrix}\bm{x}\\ \bm{y}\end{pmatrix}+\lambda^{2}_{l}\bm{x}^{\prime 2}+\alpha_{k}^{2}\bm{y}^{\prime 2}\right)\right]$
$=\exp\left(-\frac{1}{2}\bm{\omega}^{T}\begin{pmatrix}\frac{1}{\alpha_{k}^{2}}&0\\ 0&\frac{1}{\alpha_{l}^{2}}\end{pmatrix}\bm{\omega}+\begin{pmatrix}\bm{x}^{\prime}/\alpha_{k}^{2}\\ \bm{y}^{\prime}/\alpha_{l}^{2}\end{pmatrix}^{T}\bm{\omega}-\frac{\bm{x}^{\prime 2}}{2\alpha_{k}^{2}}-\frac{\bm{y}^{\prime 2}}{2\alpha_{l}^{2}}\right)$
$=\exp\left(-\frac{1}{2}\bm{\omega}^{T}A_{1}\bm{\omega}+\bm{b}_{1}^{T}\bm{\omega}+c_{1}\right)$
where $A_{1}:=\begin{pmatrix}\frac{1}{\alpha_{k}^{2}}&0\\ 0&\frac{1}{\alpha_{l}^{2}}\end{pmatrix}$, $\bm{b}_{1}:=\begin{pmatrix}\bm{x}^{\prime}/\alpha_{k}^{2}\\ \bm{y}^{\prime}/\alpha_{l}^{2}\end{pmatrix}^{T}$ and $c_{1}:=-\frac{\bm{x}^{\prime 2}}{2\alpha_{k}^{2}}-\frac{\bm{y}^{\prime 2}}{2\alpha_{l}^{2}}$. Now, we need to write $f_{X,Y}(\bm{x},y)$ in the same form:
@@
$=\bm{\omega}^{T}\begin{pmatrix}\sigma_{\epsilon}^{2}+a_{1}^{2}\sigma^{2}&-a_{1}\sigma^{2}\\ -a_{1}\sigma^{2}&\sigma^{2}\end{pmatrix}\bm{\omega}+\begin{pmatrix}2a_{0}a_{1}\sigma^{2}-2\mu\sigma_{\epsilon}^{2}\\ -2a_{0}\sigma^{2}\end{pmatrix}^{T}\bm{\omega}+a_{0}^{2}\sigma^{2}+\mu^{2}\sigma_{\epsilon}^{2},$
@@
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\exp\left(-\frac{1}{2\sigma^{2}\sigma_{\epsilon}^{2}}\left[\bm{\omega}^{T}\begin{pmatrix}\sigma_{\epsilon}^{2}+a_{1}^{2}\sigma^{2}&-a_{1}\sigma^{2}\\ -a_{1}\sigma^{2}&\sigma^{2}\end{pmatrix}\bm{\omega}+\begin{pmatrix}2a_{0}a_{1}\sigma^{2}-2\mu\sigma_{\epsilon}^{2}\\ -2a_{0}\sigma^{2}\end{pmatrix}^{T}\bm{\omega}+a_{0}^{2}\sigma^{2}+\mu^{2}\sigma_{\epsilon}^{2}\right]\right)$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\exp\left(-\frac{1}{2}\bm{\omega}^{T}\begin{pmatrix}\frac{1}{\sigma^{2}}+\frac{a_{1}^{2}}{\sigma_{\epsilon}^{2}}&-\frac{a_{1}}{\sigma_{\epsilon}^{2}}\\ -\frac{a_{1}}{\sigma_{\epsilon}^{2}}&\frac{1}{\sigma_{\epsilon}^{2}}\end{pmatrix}\bm{\omega}+\begin{pmatrix}\frac{\mu}{\sigma^{2}}-\frac{a_{0}a_{1}}{\sigma_{\epsilon}^{2}}\\ \frac{a_{0}}{\sigma_{\epsilon}^{2}}\end{pmatrix}^{T}\bm{\omega}-\frac{a_{0}^{2}}{2\sigma_{\epsilon}^{2}}-\frac{\mu^{2}}{2\sigma^{2}}\right)$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\exp\left(-\frac{1}{2}\bm{\omega}^{T}A_{2}\bm{\omega}+\bm{b}_{2}^{T}\bm{\omega}+c_{2}\right)$
@@
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{T}A_{1}\bm{\omega}+\bm{b}_{1}^{T}\bm{\omega}+c_{1}\right)\exp\left(-\frac{1}{2}\bm{\omega}^{T}A_{2}\bm{\omega}+\bm{b}_{2}^{T}\bm{\omega}+c_{2}\right)\text{d}\bm{\omega}$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{T}(A_{1}+A_{2})\bm{\omega}+(\bm{b}_{1}+\bm{b}_{2})^{T}\bm{\omega}+c_{1}+c_{2}\right)\text{d}\bm{\omega}$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{T}A\bm{\omega}+b^{T}\bm{\omega}+c\right)\text{d}\bm{\omega}$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\frac{2\pi}{|A|^{\frac{1}{2}}}\exp\left(c+\frac{1}{2}\bm{b}^{T}A^{-1}\bm{b}\right)=\frac{1}{\sqrt{\sigma^{2}\sigma_{\epsilon}^{2}|A|}}\exp\left(c+\frac{1}{2}\bm{b}^{T}A^{-1}\bm{b}\right)$

### v4

@@ B.11.1 Joint Kernel Inducing Points
Defining $\tilde{\bm{X}}:=[\tilde{\bm{x}}_{1},\tilde{\bm{x}}_{2},\dots,\tilde{\bm{x}}_{m}]^{\top}$ and $\tilde{\bm{Y}}:=[\tilde{\bm{y}}_{1},\tilde{\bm{y}}_{2},\dots,\tilde{\bm{y}}_{m}]^{\top}$, in JKIP, we optimise the following objective
@@
$=\frac{2}{m^{2}}\tilde{\bm{K}}_{m}(\tilde{\bm{x}}_{t})^{\top}\tilde{\bm{L}}_{m}(\tilde{\bm{y}}_{t})-\frac{2}{mn}\bm{K}_{n}(\tilde{\bm{x}}_{t})^{\top}\bm{L}_{n}(\tilde{\bm{y}}_{t})$ (31)
where $\tilde{\bm{K}}_{m}(\bm{x}):=[k(\bm{x},\tilde{\bm{x}}_{1}),\dots,k(\bm{x},\tilde{\bm{x}}_{m})]^{\top}$, and $\bm{K}_{n}(\bm{x}):=[k(\bm{x},\bm{x}_{1}),\dots,k(\bm{x},\bm{x}_{n})]^{\top}$, $\tilde{\bm{L}}_{m}(\bm{y}):=[l(\bm{y},\tilde{\bm{y}}_{1}),\dots,l(\bm{y},\tilde{\bm{y}}_{m})]^{\top}$, and $\bm{L}_{n}(\bm{y}):=[l(\bm{y},\bm{y}_{1}),\dots,l(\bm{y},\bm{y}_{n})]^{\top}$. This reduces the cost of evaluating the objective function by a factor of $m$, both in storage and time. Note that if using a non-stationary kernel $l(\cdot,\cdot)$, one must avoid double counting the diagonal term in Equation 30 by subtracting $\frac{1}{m}k(\tilde{\bm{x}}_{t},\tilde{\bm{x}}_{t})l(\tilde{\bm{y}}_{t},\tilde{\bm{y}}_{t})$.
@@ B.11.2 Average Conditional Kernel Inducing Points
$\underset{\text{(b)}}{=}\frac{2}{n}\tilde{\bm{L}}_{m}(\tilde{\bm{y}}_{t})^{\top}W_{\tilde{\bm{X}}\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}K_{\bm{X}\tilde{\bm{X}}}\bm{w}_{t}-\frac{2}{n}\bm{L}_{n}(\tilde{\bm{y}}_{t})^{\top}K_{\bm{X}\tilde{\bm{X}}}\bm{w}_{t}$ (32)
where (a) follows from the symmetry of the kernel functions and a simple reordering of the terms, and (b) follows from defining $\bm{w}_{t}$ to be the $t^{\text{th}}$ row of $W_{\tilde{\bm{X}}\tilde{\bm{X}}}$. Note that if using a non-stationary kernel $l(\cdot,\cdot)$ one must again subtract a correction term $\frac{1}{n}(\bm{w}_{t}^{\top}K_{\tilde{\bm{X}}\bm{X}}K_{\bm{X}\tilde{\bm{X}}}\bm{w}_{t})\cdot l(\tilde{\bm{y}}_{t},\tilde{\bm{y}}_{t})$ to avoid double counting the diagonal.
@@ Appendix C Experiment Details
$k(\bm{x},\bm{x}^{\prime}):=\exp\left(-\frac{1}{2\alpha_{k}^{2}}\|\bm{x}-\bm{x}^{\prime}\|^{2}\right),\;\;l(\bm{y},\bm{y}^{\prime}):=\exp\left(-\frac{1}{2\alpha_{l}^{2}}\|\bm{y}-\bm{y}^{\prime}\|^{2}\right)$
@@ C.1.2 Matching the Empirical Conditional Distribution - Continuous / Regression
Figure 16: Results for Heteroscedastic data with parameters set as $\bm{a}:=[3,-3,6,-6]^{\top}$, $\bm{b}:=[1,0.1,2,0.5]^{\top}$, $\bm{c}:=[-5,-2,2,5]^{\top}$, $\sigma_{1}^{2}=0.75$, and $\sigma_{2}^{2}=0.1$. RMSE is reported across a variety of test functions, as the size of the compressed set increases. For JKH (orange), JKIP (red), ACKH (blue), and ACKIP (green), we display the median performance (bold line) with the 25th-75th percentiles (shaded region) over 20 runs. The error of random sampling (black) over 500 runs is also plotted for comparison, as well as the performance of the full data (purple).
@@ C.1.3 Matching the Empirical Conditional Distribution - Discrete / Classification
Synthetic: We generate an unbalanced 4-class dataset using the multinomial logistic regression model, where conditional class probabilities are given by $\mathbb{P}(Y=0\mid X=\bm{x})=\frac{1}{1+\sum_{j=1}^{3}\exp({\bm{\beta}_{j}\cdot\bm{x}})}$ and $\mathbb{P}(Y=k\mid X=\bm{x})=\frac{\exp({\bm{\beta}_{k}\cdot\bm{x}})}{1+\sum_{j=1}^{3}\exp({\bm{\beta}_{j}\cdot\bm{x}})}$, $1\leq k\leq 3$, and $\mathbb{P}_{X}$ is a 2D Gaussian mixture model with 100 components. We assign classes using $\bm{\beta}:=\begin{bmatrix}10&8&1&45\\ 40&45&40&10\end{bmatrix}^{\top}$ and, to ensure class overlap, introduce additive noise $\epsilon\sim\mathcal{N}(0,100)$ to the exponential. This setup results in class proportions of approximately 32% (class 0), 12% (class 1), 19% (class 2), and 37% (class 3).
@@ C.3.1 Deriving the Marginal Expectation
$\mathbb{E}_{\bm{x}\sim\mathbb{P}_{X}}\left[k(\bm{x},\bm{x}^{\prime})k(\bm{x},\bm{x}^{\prime\prime})\right]=\int_{\mathcal{X}}k(\bm{x},\bm{x}^{\prime})k(\bm{x},\bm{x}^{\prime\prime})\mathbb{P}_{X}(\text{d}\bm{x})$
@@ C.3.2 Deriving the Joint Expectation
$\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}A\bm{\omega}+\bm{b}^{\top}\bm{\omega}+c\right)\text{d}\bm{\omega}$
for $\bm{\omega}:=(\bm{x},\bm{y})^{\top}$ as by completing the square, it can be shown [10] that
$\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{\top}A\bm{\omega}+\bm{b}^{\top}\bm{\omega}+c\right)\text{d}\bm{\omega}=\frac{2\pi}{|A|^{\frac{1}{2}}}\exp\left(c+\frac{1}{2}\bm{b}^{\top}A^{-1}\bm{b}\right).$
@@
$=\exp\left(-\frac{1}{2\alpha_{k}^{2}\alpha_{l}^{2}}\left[\begin{pmatrix}\bm{x}\\ \bm{y}\end{pmatrix}^{\top}\begin{pmatrix}\alpha_{l}^{2}&0\\ 0&\alpha_{k}^{2}\end{pmatrix}\begin{pmatrix}\bm{x}\\ \bm{y}\end{pmatrix}+\begin{pmatrix}-2\lambda^{2}_{l}\bm{x}^{\prime}\\ -2\alpha_{k}^{2}\bm{y}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\bm{x}\\ \bm{y}\end{pmatrix}+\lambda^{2}_{l}\bm{x}^{\prime 2}+\alpha_{k}^{2}\bm{y}^{\prime 2}\right)\right]$
$=\exp\left(-\frac{1}{2}\bm{\omega}^{\top}\begin{pmatrix}\frac{1}{\alpha_{k}^{2}}&0\\ 0&\frac{1}{\alpha_{l}^{2}}\end{pmatrix}\bm{\omega}+\begin{pmatrix}\bm{x}^{\prime}/\alpha_{k}^{2}\\ \bm{y}^{\prime}/\alpha_{l}^{2}\end{pmatrix}^{\top}\bm{\omega}-\frac{\bm{x}^{\prime 2}}{2\alpha_{k}^{2}}-\frac{\bm{y}^{\prime 2}}{2\alpha_{l}^{2}}\right)$
$=\exp\left(-\frac{1}{2}\bm{\omega}^{\top}A_{1}\bm{\omega}+\bm{b}_{1}^{\top}\bm{\omega}+c_{1}\right)$
where $A_{1}:=\begin{pmatrix}\frac{1}{\alpha_{k}^{2}}&0\\ 0&\frac{1}{\alpha_{l}^{2}}\end{pmatrix}$, $\bm{b}_{1}:=\begin{pmatrix}\bm{x}^{\prime}/\alpha_{k}^{2}\\ \bm{y}^{\prime}/\alpha_{l}^{2}\end{pmatrix}^{\top}$ and $c_{1}:=-\frac{\bm{x}^{\prime 2}}{2\alpha_{k}^{2}}-\frac{\bm{y}^{\prime 2}}{2\alpha_{l}^{2}}$. Now, we need to write $f_{X,Y}(\bm{x},y)$ in the same form:
@@
$=\bm{\omega}^{\top}\begin{pmatrix}\sigma_{\epsilon}^{2}+a_{1}^{2}\sigma^{2}&-a_{1}\sigma^{2}\\ -a_{1}\sigma^{2}&\sigma^{2}\end{pmatrix}\bm{\omega}+\begin{pmatrix}2a_{0}a_{1}\sigma^{2}-2\mu\sigma_{\epsilon}^{2}\\ -2a_{0}\sigma^{2}\end{pmatrix}^{\top}\bm{\omega}+a_{0}^{2}\sigma^{2}+\mu^{2}\sigma_{\epsilon}^{2},$
@@
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\exp\left(-\frac{1}{2\sigma^{2}\sigma_{\epsilon}^{2}}\left[\bm{\omega}^{\top}\begin{pmatrix}\sigma_{\epsilon}^{2}+a_{1}^{2}\sigma^{2}&-a_{1}\sigma^{2}\\ -a_{1}\sigma^{2}&\sigma^{2}\end{pmatrix}\bm{\omega}+\begin{pmatrix}2a_{0}a_{1}\sigma^{2}-2\mu\sigma_{\epsilon}^{2}\\ -2a_{0}\sigma^{2}\end{pmatrix}^{\top}\bm{\omega}+a_{0}^{2}\sigma^{2}+\mu^{2}\sigma_{\epsilon}^{2}\right]\right)$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\exp\left(-\frac{1}{2}\bm{\omega}^{\top}\begin{pmatrix}\frac{1}{\sigma^{2}}+\frac{a_{1}^{2}}{\sigma_{\epsilon}^{2}}&-\frac{a_{1}}{\sigma_{\epsilon}^{2}}\\ -\frac{a_{1}}{\sigma_{\epsilon}^{2}}&\frac{1}{\sigma_{\epsilon}^{2}}\end{pmatrix}\bm{\omega}+\begin{pmatrix}\frac{\mu}{\sigma^{2}}-\frac{a_{0}a_{1}}{\sigma_{\epsilon}^{2}}\\ \frac{a_{0}}{\sigma_{\epsilon}^{2}}\end{pmatrix}^{\top}\bm{\omega}-\frac{a_{0}^{2}}{2\sigma_{\epsilon}^{2}}-\frac{\mu^{2}}{2\sigma^{2}}\right)$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\exp\left(-\frac{1}{2}\bm{\omega}^{\top}A_{2}\bm{\omega}+\bm{b}_{2}^{\top}\bm{\omega}+c_{2}\right)$
@@
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{\top}A_{1}\bm{\omega}+\bm{b}_{1}^{\top}\bm{\omega}+c_{1}\right)\exp\left(-\frac{1}{2}\bm{\omega}^{\top}A_{2}\bm{\omega}+\bm{b}_{2}^{\top}\bm{\omega}+c_{2}\right)\text{d}\bm{\omega}$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{\top}(A_{1}+A_{2})\bm{\omega}+(\bm{b}_{1}+\bm{b}_{2})^{\top}\bm{\omega}+c_{1}+c_{2}\right)\text{d}\bm{\omega}$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\int_{\mathbb{R}^{2}}\exp\left(-\frac{1}{2}\bm{\omega}^{\top}A\bm{\omega}+b^{\top}\bm{\omega}+c\right)\text{d}\bm{\omega}$
$=\frac{1}{2\pi\sigma\sigma_{\epsilon}}\frac{2\pi}{|A|^{\frac{1}{2}}}\exp\left(c+\frac{1}{2}\bm{b}^{\top}A^{-1}\bm{b}\right)=\frac{1}{\sqrt{\sigma^{2}\sigma_{\epsilon}^{2}|A|}}\exp\left(c+\frac{1}{2}\bm{b}^{\top}A^{-1}\bm{b}\right)$